OpenStudy (anonymous):

Name the property of real numbers illustrated by the equation. A. Commutative Property of Multiplication B. Associative Property of Addition C. Associative Property of Multiplication D. Distributive Property i think its "B"

OpenStudy (anonymous):

@Astrophysics

OpenStudy (astrophysics):

I don't see any equation?

OpenStudy (astrophysics):

Well, here is a link that will help you with your questions of this specific content: http://www.math.com/school/subject2/lessons/S2U2L1DP.html

OpenStudy (anonymous):

3+{⌡8+11}={3+⌡8}+11

OpenStudy (anonymous):

^thats the equation^

OpenStudy (zzr0ck3r):

\(a(b+c)=ab+ac\) distribute multiplication \(a+(b+c) = (a+b) + c\) associativity of multiplication \(ab=ba\) commutative multiplication \(a+b=b+a\) commutative addition

OpenStudy (zzr0ck3r):

which one is it?

OpenStudy (anonymous):

the second one

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

awesome thank you

OpenStudy (zzr0ck3r):

that should say associativity of addition

OpenStudy (zzr0ck3r):

multiplication would be \(a(bc)=(ab)c\)

OpenStudy (anonymous):

i kew what u ment

OpenStudy (anonymous):

could u help on a few more